as ecceasea eccease in m 50 cm o
Answer #1

given

volume V = 350 cm3

= 350 x 10-6 m3  

dT = T2 - T1

= 95 - 42

= 53oC

diameter d = 5 cm

= 0.05 m

coefficient of volume expansion of water \betaw = 210 x 10-6/ oC

coefficient of volume expansion of glass \betag = 27 x 10-6/oC

dh = ?

using dh = 4 ( \betaw - \betaw ) V dT / \pi d2  

= 4 x ( 210 x 10-6 - 27 x 10-6 ) x 350 x 10-6 x 53 / 3.14 x 0.052

= 4 x 183 x 10-6 x 350 x 10-6 x 53 / 3.14 x 0.052

= 1729.757 x 10-6 m

= 1.729757 x 10-3 m

dh = 1.729757 mm